$SPAD/src/input gnarly1.input

The Axiom Team December 3, 2016

Abstract

Contents

2

— * —

draw(surface(cos(t)/(1+sin(t)**2),sin(t)*cos(t)*cos(u)/(1+sin(t)**2),
     sin(t)*cos(t)*sin(u)/(1+sin(t)**2)),t = -%pi..%pi,u = 0..%pi)

———-

References^{[1] nothing}

4